Grade 11Mechanics

A BALL DROPPED FROM A HIEGHT OF “h”m above the ground and at the same instant another ball is projected upward from the ground.The two balls meet when the upper ball fall through a distance h/3. prove that the velocities of the two balls when they meet are in the ratio 2:1

monica
8 Years agoGrade 11

1 Answer

Yash parasariya
8 Years ago
Let v1 be the velocity of the ball dropped from height h at the distance h/3 below the top.From 3rd equation of motionV1^2=u^2 +2gh Here u=0 ,h=-h/3,g=-10m/s^2V1=√2gh/3.Let v be the velocity of ball projected in ground.Time taken by this ball to reach the height 2h/3=√2h/3g (from 2nd equation of motion)From 1st equation of motion we get v=√hg/6.V1/v=2/1.